OpenStudy (anonymous):

we are starting trig subs...how do you solve integral4sin^7xcos^7xdx?

OpenStudy (frostbite):

you want to find the following \[\int\limits_{}^{}(4*\sin(x*\cos(x)^7)^7)dx\] ?

OpenStudy (anonymous):

integral of (4)(sin^7x)(cos^7x) dx

OpenStudy (anonymous):

replace sin^7x by sin(x)(sin^2x)^3 sin(x) (1-cos^2(x))^3 can you do it now ?

OpenStudy (anonymous):

your integral will become : 4sin(x)cos^7(x)(1-cos^2(x))^3 4sin(x)cos^7(x)(1-3cos^2(x)+3cos^4(x)-cos^6(x)) and this is an easy one (-sin(x) is the derivative of cos(x)..)

OpenStudy (anonymous):

are you there ?

OpenStudy (anonymous):

did i write it correctly..in that the x after the 7s are not in the power area...so it should read ^7(x) not ^7x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[4\cos(7x)\sin(7x)\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[4\cos^7(x)\sin^7(x)\]

OpenStudy (anonymous):

first or second ? which one of them

OpenStudy (anonymous):

yes on second with a dx

OpenStudy (anonymous):

so i answered you for the second one .. it is in the power area..

OpenStudy (anonymous):

=\[4\sin(x)(\sin^2(x))^3 \cos^7(x)\] =\[4\cos^7(x)\sin(x)(1-\cos^2(x))^3\] = \[4\cos^7(x)\sin(x)(1-3\cos^2(x)+3\cos^4(x)-\cos^6(x))\]

OpenStudy (anonymous):

you should be able to integrate it right now

OpenStudy (anonymous):

ok, since i wasnt sure i wanted to make sure i presented it correctly, is the above answer including the steps to solve the problem? sorry, 1st time on here not sure how to do this

OpenStudy (anonymous):

i just showed you i got to the last line .. now you have to integrate it

OpenStudy (anonymous):

great i will take it from there and try...thanks so much...much appreciated

OpenStudy (anonymous):

yw